# Confidence Calibration in Vision-Language-Action Models

### v1

@@ 4.1 Experiment Details
We perform our experiments using models derived from the OpenVLA model (Kim et al., 2025). In particular, we use models fine-tuned on 3 different task suites from the LIBERO (Liu et al., 2023) benchmark: Spatial, Object, and Goal. LIBERO is a simulation environment for language-conditioned robot manipulation tasks inspired by human activities (see Table 2 for task examples). We use 3 different fine-tuned models for each task suite: fine-tuned OpenVLA, as well as its 8-bit and 4-bit quantized versions.
Each task suite features 10 different tasks with 50 randomized initializations, for a total of 500 examples. For a given task suite, task success rate represents the proportion of the 500 trials that result in success (while task error rate represents the proportion of trials that result in failure). In order to contextualize calibration results, task success rates are reported in Table 3. To calculate ECE, we use 12 equal-mass bins and the Python package released with Kumar et al. (2019).
@@
Dataset Example Image Example Instruction
Goal [Uncaptioned image] Open the middle drawer of the cabinet.
Table 2: Task examples from the LIBERO Spatial, Object, and Goal task suites.
Model Spatial Object Goal
OpenVLA (Full) 0.828 0.880 0.758
OpenVLA (Quant-8) 0.844 0.876 0.760
OpenVLA (Quant-4) 0.822 0.850 0.712
Table 3: Task success rates across all 3 fine-tuned versions of OpenVLA and 3 different LIBERO task suites.
Figure 3: Visualization of task error rates compared against 4 different calibration error measurements for all model variants and task suites. Lower task error rates are associated with lower calibration error.
@@ 4.2 Relationship Between Calibration and Task Success
Much of the early research on calibration in deep learning models was based on perceptions of how task success (usually image classification accuracy) was related to calibration error. In particular, influential work argued that modern neural networks are poorly calibrated (Guo et al., 2017), suggesting that improved accuracy comes at the expense of calibration. This led researchers to propose training interventions to augment the cross-entropy loss with other objectives seeking improved calibration, potentially at a cost to overall accuracy (Kumar et al., 2018; Mukhoti et al., 2020). However, subsequent research found that, in fact, more accurate networks are generally better calibrated (and easier to recalibrate), and thus modifications to training procedures may not be needed (Minderer et al., 2021). Instead, techniques such as post hoc recalibration (Guo et al., 2017; Zadrozny and Elkan, 2001; Kumar et al., 2019) and ensembling (Lakshminarayanan et al., 2017; Fort and Lakshminarayanan, 2024), applied to models trained for high accuracy, are sufficient to achieve low calibration error (although calibration under distribution shift remains a significant challenge (Ovadia et al., 2019)).
To establish high-level direction for calibration research in VLAs, our first experiment focuses on this important question of how task success relates to calibration error. To study this question, in Figure 3 we plot calibration error (according to 4 different metrics) against task error rate for each of 9 combinations of 3 task suites and 3 model versions. Confidence estimates are produced using the baseline method described in Section 3.3 (a similar result under the prompt ensemble technique is shown in Appendix Figure 10).
Across all plots, a consistent trend emerges: lower task error rates are associated with lower calibration error. This mirrors findings in other areas of deep learning (e.g., as shown in image classification by Minderer et al. (2021)), and suggests that advances in task performance can naturally bring improvements in calibration. In that case, research can be focused on developing appropriate post hoc recalibration and ensembling techniques, as well as other bespoke solutions that can complement the simple and scalable learning paradigms that dominate modern deep learning.
4.3 Ensembling Confidence Over Prompts
@@
We next study the empirical effectiveness of the prompt ensemble approach described in Section 3.5. First, for the natural language instruction associated with each task, we create 20 rephrasings using GPT-4o-mini (see Appendix Table 5 for prompts used). During testing, we produce a confidence estimate conditioned on each “reprompt”, and average over these confidence scores to obtain the final model confidence. Results are reported across 3 model variants and 3 task suites, and we measure $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL, comparing to the baseline method for producing confidence estimates using the original instruction.
@@
Detailed results are recorded in Table 4 (results for the 4-bit quantized model are deferred to Appendix Table 7), while Figure 4 offers a visual summary of the high-level trends. Across all models, datasets, and metrics (36 measurements in total), the Reprompt (prompt ensemble) approach always decreases calibration error. Decreases in ECE can reach up to 40%, and average decreases are more than 20% (see Figure 11 for a visualization of percent changes across all metrics). The improvement is larger for ECE than the proper scoring rules (Brier score and NLL), where average miscalibration decreases are closer to 5-10%. This suggests that improvements in reliability (i.e., confidence matching marginal success rates) are generally greater than those in sharpness, which is expected from an ensemble technique targeted at reducing variance in confidence estimation. Overall, given its lightweight nature and strong empirical effectiveness, such data augmentation approaches seem like a promising direction for enhancing uncertainty quantification in VLAs.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier score NLL
@@
Reprompt 0.115 0.132 0.197 0.620
Quant-8 Spatial Baseline 0.070 0.093 0.138 0.488
Reprompt 0.050 0.062 0.131 0.434
Object Baseline 0.057 0.065 0.110 0.409
Reprompt 0.041 0.052 0.108 0.375
Goal Baseline 0.140 0.163 0.205 0.713
Reprompt 0.117 0.152 0.194 0.608
Table 4: Calibration error measurements for the base and 8-bit fine-tuned model versions and 3 different LIBERO task suites, for 2 different methods of confidence estimation: (1) baseline (average selected token probability) (2) Reprompt (ensembling over semantically equivalent prompts). The prompt ensemble method improves all measures across all settings.
Figure 4: Comparing calibration error measurements across 2 different methods of confidence estimation: (1) baseline (average selected token probability) (2) Reprompt (ensembling over semantically equivalent prompts).
Figure 6: Ablating the number of prompts in the prompt ensemble for the Spatial task suite. Increasing the number of prompts generally improves calibration error.
4.3.1 Robustness Ablations for Prompt Ensembles
To understand the robustness of these results, we perform multiple ablations. First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings. In addition to the prompt used to produce the results in Table 4, we try two additional prompts, recorded in Appendix Tables 5 and 6; “Prompt 1” is the original prompt, and “Prompt 2” and “Prompt 3” represent new prompts for the ablation. Note that the different prompts lead to substantially different rephrasings, as Prompts 2 and 3 mandate the retention of the words “pick” and “place”, while as Table 1 shows, Prompt 1 leads to these words being replaced. The ablation is run across all 3 task suites using fine-tuned OpenVLA, and we measure $\text{ECE}_{1}$ and NLL. Results are shown in Figure 5. Improvements in calibration error from the Reprompt method are robust to different rephrasing prompts, with all ensemble variants producing lower error than the baseline method across both metrics and all 3 task suites.
Second, we consider the effect of the number of prompts used in the prompt ensemble. Given a sound ensembling approach, we would expect to see improvement as more prompts are added to the ensemble (up to some point). To study this question, we record results using $k\in\{1,5,10,20\}$ different rephrasings (randomly chosen for 1000 trials when $k<20$). Results for the Spatial task suite are shown in Figure 6, and additional results for the Goal and Object task suites are in Appendix Figure 12. We see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
@@ 4.4 Calibration Over Task Time
To test this intuition we measure calibration across 500 test trials for 100 different levels of task completion ($\{0,1,\dots,99\}\%$), where task completion is calculated as the current timestep index $t$ divided by the total number of timesteps in the task episode (and multiplied by 100). Our goal is to understand whether the quality of confidence estimates improves as the robot progresses in its task. At each measured timestep, we compute $\text{ECE}_{1}$ and Brier score using the baseline method for producing confidence estimates. Because a downstream safety monitor might consider the history of recent estimates (beyond the current one), we evaluate three aggregation rules for reporting a confidence estimate at each timestep:
@@
To better understand the calibration error results, we also visualize how average confidence estimates evolve over the task horizon, separated by successful and failed trials. Finally, we plot reliability diagrams for the timesteps corresponding to $\{0,50,75,99\}\%$ completion. Reliability diagrams (Guo et al., 2017) offer a visualization of expected calibration error, using a similar binning strategy. They show confidence against accuracy for each bin, where a well-calibrated system lies on the $x=y$ line. Results for the Spatial task suite with the fine-tuned OpenVLA model are shown in Figure 7. Additional results for Spatial with the Quant-8 model (Figure 13), as well as the Object and Goal suites with the full precision (Figures 14 and 16) and Quant-8 models (Figures 15 and 17), are presented in Appendix B.3.
@@
Figure 7: Empirical study of calibration error across the task time horizon for the Spatial task suite and the full precision fine-tuned OpenVLA model. In the top row, the left two plots characterize how calibration error evolves with task progress, while the right two plots show the average confidence estimates by task time, grouped by successful and failed trials. The bottom two rows offer a sample of the reliability diagrams produced by different methods for aggregating confidence estimates across time. Overall, these results illustrate that calibration can improve as the task progresses and more information is gathered, suggesting opportunities for context-aware uncertainty interventions.
@@
Focusing first on results using the confidence estimate from the current step, a clear pattern emerges. Across all 6 settings (3 task suites each with 2 model variants) and both metrics, calibration improves sharply from 0% to approximately 50% completion, then plateaus or deteriorates back towards the original level (left-most and left-middle plots in the top row of the results figures). Beyond lower scores on calibration error, our additional plots characterize the improved probabilistic nature of the confidence estimates towards the middle of the task horizon. In the right-middle and right-most plots in the top row (titled “Successful Trials” and “Failed Trials”), the gap between the average confidence on successful vs. failed trials tends to be greatest around 50% task completion. Additionally, considering the top row of reliability diagrams, we can see that the confidence estimates using the current step become far more reliable around this point: error for many bins is large at the beginning, then the difference between confidence and accuracy for most bins becomes smaller until the task is roughly 75% complete. Near task completion, the gap between confidence on successful vs. failed trials closes again, and reliability suffers as a result.
These results support our hypothesis regarding the utility of gathering more information before expressing a confidence estimate, even using the simple baseline method. Early miscalibration might be expected: with only a high-level image frame and a textual instruction, the policy lacks much information about how hard the task might be. By mid-trajectory, it might observe how its gripper and the scene interact, analogous to a “test grip” on the wine glass, so its probability estimates may better match reality. Near the goal state, however, calibration error is generally worse, for instance because of differences in the action dimensions being emphasized in each phase of the task. These experiments suggest a practical recipe: let the robot execute a certified-safe prefix of the trajectory, and then assess its confidence and intervene if necessary. Such horizon-aware monitoring balances safety (no contact made during the prefix) with the improved reliability in confidence estimates that comes from extra perceptual context.
@@
Beyond the baseline strategy of estimating confidence based only on the current timestep, we next observe the effects of other simple strategies for aggregating across timesteps. Specifically, we consider a sliding window average over 5 single-step estimates, and the average of all observed single-step estimates. Across all model/task suite combinations, the sliding window method is particularly effective at improving the Brier score around the middle of the task horizon, although ECE is not always improved. With respect to averaging all single-step estimates so far, we observe that Brier score improves throughout the task, but ECE generally gets worse. The underlying behavior driving these changes can be observed in the bottom row of reliability diagrams: the “Avg. All” method is fairly successful at sorting the failed examples to have relatively lower confidence, but also reduces variance in the confidence estimates such that they are highly overconfident in this range. Thus, while this approach fares poorly according to ECE, it is actually promising in the sense that it enables more effective discrimination between successful and failed trials.
@@ 4.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
To illustrate how context-aware confidence monitoring could work in practice, we apply it to a representative pick and place task from the Goal suite. The task is to “put the wine bottle on the rack”, a case where the robot should be relatively conservative to avoid breaking glass. We consider a naive approach to context-aware monitoring, proposing to halt task performance when both:
  1. 1.
    The confidence level falls below a threshold set to the 10% quantile of confidence estimates for that point in the task horizon across all task trials (based on percent completion).
  2. 2.
    The robot is within a few inches of contacting an object, or already has contacted an object.
Since building a system to detect proximity to objects is beyond the scope of this work, we focus on a qualitative demonstration of this idea. For each example under examination, we plot current confidence for $\{0,20,40,60,80,98\}\%$ completion, as well as the corresponding 10% quantile risk threshold and an image of the robot environment at that time.
Some particularly illustrative examples are shown in Figure 8. First, we consider the episode shown across the top row (Episode 29 of the 50 in the Goal task suite). Here, confidence begins high, but falls until it is slightly below the halting threshold at 20% task completion, coinciding with the difficult subtask of gripping the rounded glass bottle. Allowed to continue, the model is able to grip the bottle, and confidence rebounds as it approaches the wine rack. However, confidence once again falls quickly when it begins to set the bottle down, possibly because the difficult grip put the bottle in an unfavorable position for placement. In the end, the robot fails to place the bottle securely on the rack, with potentially negative real-world consequences. Given the opportunity, a safety intervention could have been performed at multiple points before this incident, either to avoid grasping the bottle in the first place or to reset the bottle safely on the table instead of trying a failed placement.
@@
In both Episode 46 (middle row) and Episode 49 (bottom row), confidence begins low and below the threshold. However, given the knowledge that calibration improves throughout the task horizon, we may prefer to allow the robot to proceed with the task as long as it is not contacting or nearly contacting any objects. Comparing these examples shows the potential of such an approach. In Episode 46, although confidence increases throughout the beginning of the task, it dives sharply after 40% task completion and remains low throughout the rest of the task, as the wine bottle is ultimately knocked down and the model is unable to recover. Given our proposed strategy, the task might have been halted before the wine bottle was contacted (at 40% completion), and could, for example, have been deferred to a human. On the other hand, in Episode 49 confidence estimates remain above the threshold for the rest of the trial, and the robot is successful. This episode also highlights the usefulness of an adaptive confidence threshold, given that the confidence estimate at 98% task completion would have fallen below the threshold at other timesteps.
More such qualitative examples are provided in Appendix Figure 18. These include cases where the strategy succeeds and others where it fails (e.g., confidence falls below the threshold while grasping in an ultimately successful trial).

### v2

@@ 4.1 Experiment Details
We perform our experiments using 4 different VLA variants: OpenVLA, MolmoAct (lee2025molmoactactionreasoningmodels), UniVLA (bu2025univlalearningacttaskcentric), and NORA (hung2025norasmallopensourcedgeneralist). We choose these models because they each predict tokens in a manner that allows for a natural probabilistic interpretation. [^2] Descriptions of confidence estimation with MolmoAct, UniVLA, and NORA are found in Appendix A.1. All models are fine-tuned on 4 different task suites from the LIBERO (liu2023liberobenchmarkingknowledgetransfer) benchmark: Spatial, Object, Goal, and 10. LIBERO is a simulation environment for language-conditioned robot manipulation tasks inspired by human activities (see Table 3 for task examples). We also include results for the 8-bit and 4-bit quantized versions of OpenVLA fine-tuned on Spatial, Object, and Goal, for a total of 22 model/task suite combinations.
Each task suite features 10 different tasks with 50 randomized initializations, for a total of 500 examples. For a given task suite, task success rate represents the proportion of the 500 trials that result in success (while task error rate represents the proportion of trials that result in failure). To calculate ECE, we use 12 equal-mass bins and the Python package released with kumar2020verifieduncertaintycalibration.
@@ 4.2 Relationship Between Calibration and Task Success
Much of the early research on calibration in deep learning models was based on perceptions of how task success (usually image classification accuracy) was related to calibration error. In particular, influential work argued that modern neural networks are poorly calibrated (guo2017calibration), suggesting that improved accuracy comes at the expense of calibration. This led researchers to propose training interventions to augment the cross-entropy loss with other objectives seeking improved calibration, potentially at a cost to overall accuracy (pmlr-v80-kumar18a; mukhoti2020calibratingdeepneuralnetworks). However, subsequent research found that, in fact, more accurate networks are generally better calibrated (and easier to recalibrate), and thus modifications to training procedures may not be needed (minderer2021revisiting). Instead, techniques such as post hoc recalibration (guo2017calibration; zadrozny2001obtaining; kumar2020verifieduncertaintycalibration) and ensembling (lakshminarayanan2017simple; fort2024ensembleeverywheremultiscaleaggregation), applied to models trained for high accuracy, are sufficient to achieve low calibration error (although calibration under distribution shift remains a significant challenge (ovadia2019trust)).
To establish high-level direction for calibration research in VLAs, our first experiment focuses on this important question of how task success relates to calibration error (according to $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL) across the 22 model/task suite combinations described above. Results are visualized in Figure 3, and also reported in Table 6. All models exhibit a roughly monotonic relationship between task error and the discriminative metrics (Brier score and NLL), but they differ on ECE: OpenVLA and MolmoAct tend to achieve lower ECE when task error is low, whereas UniVLA and NORA do not show as clean a trend. One possible explanation is that in OpenVLA and MolmoAct, the cross-entropy loss (a proper scoring rule) more directly supervises discretized action dimension tokens, so bin-averaged confidence better tracks per-task success. By contrast, the latent or compressed action representations and auxiliary objectives in UniVLA and NORA may introduce a more significant mismatch between token probabilities and success. These patterns are only suggestive, but they point to architectural complexity in modern VLAs as a potential source of unfavorable calibration behavior. Future work with controlled ablations is needed to fully understand which design choices drive the observed behavior.
Figure 3: Visualization of task error rates compared against 4 different calibration error measurements for 4 VLA variants (OpenVLA, MolmoAct, UniVLA, and NORA) and 4 LIBERO task suites (Spatial, Object, Goal, 10), as well as OpenVLA 8- and 4-bit versions on Spatial, Object, and Goal. All models exhibit a roughly monotonic relationship between task error and the discriminative measures (Brier score and NLL). ECE shows differences between model families, potentially due to architecture and objective differences.
4.3 Ensembling Confidence Across Prompts
@@
We next study the empirical effectiveness of the prompt ensemble approach described in Section 3.5. First, for the natural language instruction associated with each task, we create 20 rephrasings using GPT-4o-mini (see Appendix Table 4 for the prompts used). During testing, we produce a confidence estimate conditioned on each “reprompt”, and average over these confidence scores to obtain the final model confidence. We evaluate OpenVLA and its quantized variants across 3 task suites, as well as UniVLA and NORA on the Spatial suite. We measure $\text{ECE}_{1}$, $\text{ECE}_{2}$, Brier score, and NLL, comparing to the baseline method for producing confidence estimates using the original instruction.
@@
Detailed results are recorded in Table 2 (results for the 8-bit and 4-bit quantized models are deferred to Appendix Table 7). Across all models and task suites, the Reprompt (prompt ensemble) approach always decreases calibration error according to $\text{ECE}_{1}$, $\text{ECE}_{2}$, and NLL, and never performs worse according to any of the metrics. Decreases in ECE can reach up to 40%, and average reductions are roughly 20% (see Figure 7 for a visualization of percent changes across all metrics). The improvement is larger for ECE than the proper scoring rules (Brier score and NLL), where average miscalibration decreases are closer to 5-10%. This suggests that improvements in reliability (i.e., confidence matching marginal success rates) are generally greater than those in sharpness, which is expected from an ensemble technique targeted at reducing variance in confidence estimation. Overall, given its lightweight nature and strong empirical effectiveness, such data augmentation approaches seem like a promising direction for enhancing uncertainty quantification in VLAs.
Model Dataset Method $\text{ECE}_{1}$ $\text{ECE}_{2}$ Brier NLL
@@
Goal Reprompt 0.115 0.132 0.197 0.620
UniVLA Spatial Baseline 0.162 0.193 0.064 0.282
Spatial Reprompt 0.157 0.192 0.064 0.262
NORA Spatial Baseline 0.099 0.115 0.105 0.386
Spatial Reprompt 0.092 0.112 0.105 0.376
Table 2: Calibration error measurements for 2 different methods of confidence estimation: (1) baseline (average selected token probability); (2) Reprompt (ensembling over semantically equivalent prompts). Prompt ensembling consistently improves all measures, and never leads to worse calibration.
To understand the robustness of these results, we perform multiple ablations; because of space considerations, we defer full details and results to Appendix B.3. First, we consider the effect of changing the prompt given to GPT-4o-mini for producing the 20 instruction rephrasings, finding that improvements in calibration error from the Reprompt method are robust to different rephrasing prompts. Second, we consider the effect of the number of prompts used in the prompt ensemble, and see that the algorithm behaves favorably, where calibration error generally decreases as more instructions are included in the ensemble.
@@ 4.4 Calibration Over Task Time
To test this intuition, we measure calibration across 500 test trials for 100 different levels of task completion ($\{0,1,\dots,99\}\%$), where task completion is calculated as the current timestep index $t$ divided by the total number of timesteps in the task episode (and multiplied by 100). Results for each level of task completion are averaged across the 500 trials, to study whether high-level confidence and calibration trends might occur. In this section, we focus on 6 model/task suite combinations (OpenVLA and its 8-bit version, each applied to Spatial, Object, and Goal suites), to understand how any observations might generalize. Our goal is to examine whether the quality of confidence estimates changes as the robot progresses in its task. At each measured timestep, we compute $\text{ECE}_{1}$ and Brier score using the baseline method for producing confidence estimates. Because a downstream safety monitor might consider the history of recent estimates (beyond the current one), we evaluate three aggregation rules for reporting a confidence estimate at each timestep:
@@
Beyond calibration error, we also visualize how average confidence estimates evolve over the task horizon, separated by successful and failed trials. Finally, we plot reliability diagrams for the timesteps corresponding to $\{0,50,75,99\}\%$ completion. Reliability diagrams (guo2017calibration) offer a visualization of expected calibration error, using a similar binning strategy. They show confidence against accuracy for each bin, where a well-calibrated system lies on the $x=y$ line. Results for the Spatial task suite with the fine-tuned OpenVLA model are shown in Figure 4. Additional results for Spatial with the Quant-8 model (Figure 10), as well as the Object and Goal suites with the full precision (Figures 11, 13) and Quant-8 models (Figures 12, 14), are presented in Appendix B.4.
@@
Figure 4: Empirical study of calibration error across the task time horizon. In the top row (a), the left two plots show how calibration evolves with task progress, while the right two plots show the average confidence by task time, grouped by successful and failed trials. The bottom two rows (b, c) offer a sample of the reliability diagrams produced by different methods for aggregating confidence estimates across time. Overall, these results illustrate that calibration can improve as the task progresses and more information is gathered, suggesting opportunities for context-aware uncertainty interventions.
@@
Focusing first on results using the confidence estimate from the current step, a clear pattern emerges. As shown in Figure 4 and Figures 10-14, across all 6 settings and both metrics calibration improves sharply from 0% to approximately 50% completion, then plateaus or deteriorates back towards the original level (left-most and left-middle plots in section (a) of the results figures). Beyond lower scores on calibration error, our additional plots characterize the improved probabilistic nature of the confidence estimates towards the middle of the task horizon. In the right-middle and right-most plots in the top section (a) (titled “Successful Trials” and “Failed Trials”), the gap between the average confidence on successful vs. failed trials tends to be greatest around 50% task completion. Additionally, considering the top row of reliability diagrams (section (b)), we can see that the confidence estimates using the current step become far more reliable around this point: error for many bins is large at the beginning, then the difference between confidence and accuracy for most bins becomes smaller until the task is roughly 75% complete. Near task completion, the gap between confidence on successful vs. failed trials closes again, and reliability suffers as a result. These results suggest a practical recipe: let the robot execute a certified-safe prefix of the trajectory, and then assess its confidence and intervene if necessary. Such horizon-aware monitoring balances safety (no contact made during the prefix) with the improved reliability in confidence estimates that comes from extra perceptual context.
@@
We next observe the effects of other simple strategies for aggregating confidence estimates across timesteps. As described above, alongside the current step confidence, we also consider a sliding window average over 5 single-step estimates and the average of all observed single-step estimates. Across all model/task suite combinations, the sliding window method is particularly effective at improving the Brier score around the middle of the task horizon, although ECE is not always improved. With respect to averaging all single-step estimates so far, we observe that Brier score improves throughout the task, but ECE generally gets worse. The underlying behavior driving these changes can be observed in the bottom row of reliability diagrams (marked (c)): the “Avg. All” method is successful at assigning relatively lower confidence to failed examples, but also reduces variance in the estimates such that they are highly overconfident in these bins (i.e., for the leftmost bin at 99% completion, confidence is still high, but accuracy is near zero). Thus, while this approach fares poorly according to ECE, it is actually promising in the sense that it enables more effective discrimination between successful and failed trials.
Finally, to further probe the generality of these results, we repeat these experiments using the UniVLA model, presented in Appendix Figures 15, 16, and 17. We once again find that calibration improves after making some task progress before deteriorating again, suggesting opportunities for context- and risk-aware applications of confidence quantification.
@@ 4.4.1 Qualitative Examples of Context-Aware Confidence Monitoring
To illustrate how context-aware confidence monitoring could work in practice, we apply it to a representative pick and place task from the Goal suite. The task is to “put the wine bottle on the rack”, a case where the robot system should be relatively conservative to avoid breaking glass. We consider a naive approach to context-aware monitoring, proposing to halt task performance when both: (1) the confidence level falls below a threshold set to the 10% quantile of confidence estimates for that point in the task horizon across all task trials (based on percent completion); (2) the robot is within a few inches of contacting an object, or already has contacted an object. Since building a system to detect proximity to objects is beyond the scope of this work, we focus on a qualitative demonstration of this idea. For each example under examination, we plot current confidence for $\{0,20,40,60,80,98\}\%$ completion, as well as the corresponding 10% quantile risk threshold and an image of the robot environment at that time.
@@
Some particularly illustrative examples are shown in Figure 5. First, we consider the episode shown across the top row (marked (a), Episode 29 of the 50 in the Goal task suite). Here, confidence begins high, but falls until it is slightly below the halting threshold at 20% task completion, coinciding with the difficult subtask of gripping the rounded glass bottle. Allowed to continue, the model is able to grip the bottle, and confidence rebounds as it approaches the wine rack. However, confidence once again falls quickly when it begins to set the bottle down, possibly because the difficult grip put the bottle in an unfavorable position for placement. In the end, the robot fails to place the bottle securely on the rack, with potentially negative real-world consequences. Given the opportunity, a safety intervention could have been performed at multiple points before this incident, either to avoid grasping the bottle in the first place or to reset the bottle safely on the table instead of trying a failed placement.
In both Episode 46 (marked (b)) and Episode 49 (marked (c)), confidence begins below the threshold. However, given the knowledge that calibration improves throughout the task horizon, we may prefer to allow the robot to proceed with the task as long as it is not too near any objects. Comparing these examples shows the potential of such an approach. In Episode 46, although confidence increases throughout the beginning of the task, it dives sharply after 40% task completion and remains low throughout the rest of the task, as the wine bottle is ultimately knocked down and the model is unable to recover. Given our proposed strategy, the task might have been halted before the wine bottle was contacted (at 40% completion), and could, for example, have been deferred to a human. On the other hand, in Episode 49 confidence estimates remain above the threshold for the rest of the trial, and the robot is successful. This episode also highlights the usefulness of an adaptive confidence threshold, given that the confidence estimate at 98% task completion would have fallen below the threshold at other timesteps.
More such qualitative examples are provided in Appendix Figure 18. These include both cases where the strategy succeeds and others where it fails (e.g., in Episode 1 confidence falls below the threshold while grasping in an ultimately successful trial).